# Provable Mixed-Noise Learning with Flow-Matching

### v1

@@ 5.1 MNIST Denoising
We apply our flow matching noise estimation algorithm to MNIST denoising. For this we choose $a_{true}=0.1$ and $b_{true}=0.3$ and rescale the MNIST images to [1,1][-1,1]. We generate 16 noisy measurements, as can be seen on the left of Fig. 2. Then we iterate between the flow matching algorithm for the current estimates of $(a,b)$ and then estimate $(a,b)$ based on the measurements. For the final flow matching model, one can see the posterior (aka clean) samples of the flow matching model on the right of Fig. 2. We show the convergence of a and b over the epochs in Fig. 3. Here, $a$ decreases towards the true value and b increases towards the true value. The initial estimation was both for a and b at $\frac{1}{2}$.
@@ 5.2 Reaction-Diffusion PDE
We approximate the forward operator based on the PDE bench dataset as provided in [28] using a Fourier Neural Operator [17, 16]. Further, our velocity field in the flow matching framework is also parameterized by an FNO, see also [12].
@@ 6 Conclusions
Still, as outlined in remark 10, a fruitful future research direction is to sharpen the bounds, so that the convergence condition $\gamma<c$ can be guaranteed.

### v2

@@ 5.1 MNIST Denoising
We apply our flow matching noise estimation algorithm to MNIST [21] denoising. For this we choose $a_{true}=0.1$ and $b_{true}=0.3$ and rescale the MNIST images to [1,1][-1,1]. We generate 16 noisy measurements, as can be seen on the left of Fig. 2. Then we iterate between the flow matching algorithm for the current estimates of $(a,b)$ and then estimate $(a,b)$ based on the measurements. For the final flow matching model, one can see the posterior (aka clean) samples of the flow matching model on the right of Fig. 2. We show the convergence of a and b over the epochs in Fig. 3. Here, $a$ decreases towards the true value and b increases towards the true value. The initial estimation was both for a and b at $\frac{1}{2}$.
@@ 5.2 Reaction-Diffusion PDE
We approximate the forward operator based on the PDE bench dataset as provided in [33] using a Fourier Neural Operator [22, 19]. Further, our velocity field in the flow matching framework is also parameterized by an FNO, see also [15]. By this, we mean that we concatenate the time input with the images in a channel and use a standard FNO.
The choice of the neural operator does not correspond to an infinite dimensional framing of this inverse problem, but rather that the FNO has shown to be effective in modelling PDE solutions [22, 33].
@@ 6 Conclusions
Still, it is unclear whether our EM convergence theory applies to noise models with compact support, and whether it can be shown that our probabilistic FOS translates into EM convergence under the conditions on $\gamma$ and $c$.